OpenStudy (anonymous):

Write the measure in radians. Express the answer in terms of p. 320º

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (amistre64):

multiply your degrees by pi/180 to convert to radians

OpenStudy (amistre64):

\[\frac{320}{1}*\frac{\pi}{180}=\frac{320\pi}{180}\]

OpenStudy (amistre64):

reduce as wanted

OpenStudy (anonymous):

THANK YOU.

OpenStudy (amistre64):

youre welcome :)